Sps ls1 planning meeting no 39 25 06 14
Download
1 / 41

SPS LS1 Planning meeting (No.39) 25/06/14 - PowerPoint PPT Presentation


  • 57 Views
  • Uploaded on

SPS LS1 Planning meeting (No.39) 25/06/14. David Mcfarlane. Topics. Lifts and Monte-charges . GE/ASE Access tests. Cleaning campaign dates. EPC tests of the Auxiliary Converters. The end of LS1. Access after the 27 th June. Planning. Rubbish. E.C.Rs. Important Information .

loader
I am the owner, or an agent authorized to act on behalf of the owner, of the copyrighted work described.
capcha
Download Presentation

PowerPoint Slideshow about ' SPS LS1 Planning meeting (No.39) 25/06/14' - penn


An Image/Link below is provided (as is) to download presentation

Download Policy: Content on the Website is provided to you AS IS for your information and personal use and may not be sold / licensed / shared on other websites without getting consent from its author.While downloading, if for some reason you are not able to download a presentation, the publisher may have deleted the file from their server.


- - - - - - - - - - - - - - - - - - - - - - - - - - E N D - - - - - - - - - - - - - - - - - - - - - - - - - -
Presentation Transcript
Sps ls1 planning meeting no 39 25 06 14

SPS LS1Planning meeting (No.39)25/06/14

David Mcfarlane

[email protected]


Topics
Topics

  • Lifts and Monte-charges.

  • GE/ASE Access tests.

  • Cleaning campaign dates.

  • EPC tests of the Auxiliary Converters.

  • The end of LS1.

  • Access after the 27th June.

  • Planning.

  • Rubbish.

  • E.C.Rs.

  • Important Information.

  • A.O.B.

[email protected]





Epc tests of the auxiliary converters
EPC tests of the Auxiliary Converters

  • With this planning, OP can start the commissioning of the Software with a 1 week delay from our activities: e.g, OP tests for BA2 could start on 9th June.   

  • Whatweneed:

  • The magnet interlocks shall be tested in May by MPE (R.Mompo/P.Dahlen): support from EPC is requested during this period,

  • Services shall be present in each BA in time: water (magnet + surface building), no consignation (except BA1 until end of June),

  • A general signalisation that some tests are ongoing: coordination EPC/MSC,

  • EIS (https://edms.cern.ch/document/991272): from past experiences, this is a difficult point for us during our tests, especially for TI2/TI8. I propose that when the planning is agreed on, Marc Tavlet will be contacted by EPC.

[email protected]



Access after the 27 th june
Access after the 27th June.

  • ALL LS1-SPS IMPACTs will end on the 27th June!

  • Any accesses into any area of the SPS after this date will require a NEW IMPACT.

  • You must enter the correct “intervention period” depending on where you will be accessing.

    • LS1-HWT-SPS(valid from 28th June to 31st August)

      • Access into the SPS during Hardware commissioning phase

    • LS1-BA80-SPS(valid from now to 30th September)

      • Allowing continued access into BA80

    • TAG41- AWAKE (Valid from now until October 2016)

      • Allowing continued access into TAG41

[email protected]


LS1-BA80-SPS

(valid from now to 30th September)

LS1-HWT-SPS

(valid from 28th June to 31st August)

TAG41- AWAKE

(Valid from now until October 2016)

[email protected]


Access during july august
Access during July/August

  • Access will once again be controlled by the CCC.

  • There will be informal morning meetings everyday in the CCC to discuss if and when any accesses can be made.

  • There is a meeting in the CCC today at 16:30 to finalize these details.

[email protected]


Ba1 works
BA1 works

  • Reinstallation of the beam line and equipment in LSS1

    • https://edms.cern.ch/document/1293491/6.0/TAB3

[email protected]






What a load of rubbish
What a load of RUBBISH!!

  • Its very important to ensure that you leave the work site clean and tidy.

  • REMOVE ALL OF YOUR RUBBISH FROM THE TUNNEL NOW!!!!!!!!!

[email protected]

















Engineering change requests
Engineering Change Requests

  • Any and all new equipment being installed in the SPS requires an ECR BEFORE it can be installed. It must also be approved by planning and coordination.

  • Templates and Guidelines for Engineering Change Requests:

    • EDMS: 1271880

  • SPS ECR template (in DFS):

    • https://dfsweb.web.cern.ch/dfsweb/Services/DFS/DFSBrowser.aspx/Applications/CERN/Workgroup%20Templates/SPS/ECR-Template-SPS.dotx

[email protected]


Current ecr s for ls1
Current ECR’s for LS1

https://edms.cern.ch/nav/P:CERN-0000076703:V0/P:CERN-0000096067:V0/TAB3

  • Removal of AEWB Pick-up from position 31932 to 31502

    • (EDMS No. 1245361) UNDER IMPLEMENTATION

  • Removal of a BPWA Pick-up from position 31931 to 31101

    • (EDMS No. 1245393) UNDER IMPLEMENTATION

  • Installation of a new Wideband Kickers system in position 32173

    • (EDMS No. 1245400)UNDER IMPLEMENTATION

  • Installation of a new HBW Damper Pick-up in place of BPWA 31931 & AEWB 31932

    • (EDMS No. 1245401)RELEASED

  • Replacement of 4 old BPA and 2 BPD by 6 BPCE in LSS1 & LSS2

    • (EDMS No. 1246559)UNDER IMPLEMENTATION

  • Reinstallation of pump and gauges (Pirani + Penning) in 11993 (vacuum sector 150)

    • (EDMS No. 1249305)RELEASED

  • Modification LSS1 sectorization (new VVSF + displacement VVSB 11772)

    • (EDMS No. 1249307 v.2) RELEASED

  • Installation of 3 BCT in LSS5 for North Area protection during Ions operations in SPS.

    • (EDMS No. 1075945) UNDER IMPLEMENTATION

[email protected]


Current ecr s for ls11
Current ECR’s for LS1

https://edms.cern.ch/nav/P:CERN-0000076703:V0/P:CERN-0000096067:V0/TAB3

  • Transfer of the Fast Beam Current Transformers (BCTFR.31451) for the SPS circulating beam from LSS3 to LSS5.

    • (EDMS No. 1221341) RELEASED

  • SPS Beam Position Monitoring Prototyping Setup in LSS4.

    • (EDMS No. 1099630)RELEASED

  • Installation de 4 moniteurs de position (2 BPH, 2 BPV), dédiés au système Damper.

    • (EDMS No. 1251693)RELEASED

  • Installation de 2 nouveaux moniteurs de position de type BPCR en LSS2 au 21459 et 22172.

    • (EDMS No. 1251695) UNDER IMPLEMENTATION

  • Displacement of Door PPG2145 separating the LHC and the SPS in the TI2 tunnel.

    • (EDMS No. 1253832) UNDER IMPLEMENTATION

  • Interlocking of Injection Tunnel upstream areas with LHC powering.

    • (EDMS No. 1257949)RELEASED

  • Move of BPH.31998 to BPH.42173.

    • (EDMS No. 1272718) RELEASED

  • Optical Fibre Racks.

    • (EDMS No. 1273766)UNDER IMPLEMENTATION

[email protected]


Current ecr s for ls12
Current ECR’s for LS1

https://edms.cern.ch/nav/P:CERN-0000076703:V0/P:CERN-0000096067:V0/TAB3

  • Removal of old fixed TMP vacuum groups in 11655-11691.

    • (EDMS No.1273232)UNDER IMPLEMENTATION

  • BYBPM racks under the SPS dipole magnets as part of the Upgrade of Beam Position Monitors electronics (MOPOS).

    • (EDMS No. 1282651) UNDER IMPLEMENTATION

  • Installation de Coffrets d'Alimentation en LSS1 et TT10.

    • (EDMS No. 1282860) UNDER IMPLEMENTATION

  • Installation d'un rack de jonction RX11695=TS1 au BA1, afin de limiter les longueurs des câbles à changer lors d'un futur remplacement.

    • (EDMS No. 1288607)UNDER IMPLEMENTATION

  • Installation de nouveaux chemins de câbles en LSS1+ du SPS.

    • (EDMS No. 1291889 v.2) RELEASED

  • TPSN replacement in LSS2 of SPS.

    • (EDMS No. 1302672 v.1) UNDER IMPLEMENTATION

  • Replacement of pumps in 3 tanks MKP in LSS1 of the SPS .

    • (EDMS No. 1328248 v. 0.1)RELEASED

  • Suppression of VVFA 11899 in LSS1 of SPS.

    • (EDMS No. 1328249v. 0.1) RELEASED

[email protected]


Current ecr s for ls13
Current ECR’s for LS1

https://edms.cern.ch/nav/P:CERN-0000076703:V0/P:CERN-0000096067:V0/TAB3

  • New shielding wall behind TIDV 11892 in SPS .

    • (EDMS No. 1331621v. 0.1)RELEASED

  • Modification of Numbering of SPS Civil Works.

    • (EDMS No. 1346653v. 0.1)UNDER IMPLEMENTATION

  • New Cable Tray in LSS5- of SPS.

    • (EDMS No. 1346655v. 0.1)UNDER IMPLEMENTATION

  • Helium 2K Pumping Units Installation in the Alcove Region of SPS Point 4 (BA4).

    • (EDMS No. 1354826v. 1.0)UNDER IMPLEMENTATION

  • Installation of a wire scanner BWSRE in LSS5 of the SPS.

    • (EDMS No. 1356736v. 0.1)UNDER IMPLEMENTATION

  • Increase of the MDHD 11832 steering strength for the LIU-SPS Q20 optics.

    • (EDMS No. 1359692v. 0.1)RELEASED

  • Modification of 2 Shielding Towers in LSS5 of the SPS.

    • (EDMS No. 1360306v. 0.1)UNDER IMPLEMENTATION

  • Upgrade of the UA9 equipment during LS1.

    • (EDMS No. 1365265v. 1.0) UNDER IMPLEMENTATION

[email protected]


Current ecr s for ls14
Current ECR’s for LS1

https://edms.cern.ch/nav/P:CERN-0000076703:V0/P:CERN-0000096067:V0/TAB3

  • Removal of a BCT in TT10 during LS1.

    • (EDMS No. 1368378v. 1.0) RELEASED

  • Modification of the magnetic bumps for safe operation of the BIPM in SPS LSS5.

    • (EDMS No. 1377890v. 1.0)UNDER IMPLEMENTATION

  • New in situ SEY measurement system in LSS5 of the SPS.

    • (EDMS No. 1380468v. 1.0)IN WORK

  • New HBW Damper Kickers.

    • (EDMS No. 1391500v. 0.1)IN WORK

[email protected]


Important information
Important information

  • Useful Contact Numbers

    • SPS Coordination Team

      • David MCFARLANE- 164247

      • Franck BAIS – 165467

      • James RIDEWOOD - 160590

    • Control Room

      • SPS CCC – 77500

      • TI CCC – 72201

      • CSA - 62511

    • RP

      • Christophe Tromel – 163199

      • Angelito Herve – 163168

      • Nadine Conan (BA1) – 160641

      • Elodie Aubert - 169097

      • Kurt Weiss (BA1) -160759

    • Coordination Safety

      • Emmanuel Paulat – 163870

      • John Etheridge - 164647

[email protected]


A o b
A.O.B.

  • There are currently 58 IMPACTs that are “LATE”. If you have completed your work then you need to close the related IMPACT.

[email protected]



Sps coordination website
SPS coordination website

http://sps.web.cern.ch/SPS/

http://cern.ch/LS1planning

\\cernhomeJ.cern.ch\J\jocksoft\Public\LS1

Next meeting

To Be Annnounced

[email protected]


ad